OpenStudy (anonymous):

Please help. I've worked on this problem for two days :( What are the vetices of the ellipse equation: x^2 + 10x +4y^2-56y=-205

OpenStudy (anonymous):

\[(9,7),(-1,7)\]

OpenStudy (anonymous):

I think it should be (-9,7), (-1, 7)

OpenStudy (anonymous):

yes you are right. i forgot the minus sign

OpenStudy (anonymous):

first complete the square to get \[(x+5)^2+4 (y-7)^2= 16\] then \[\frac{(x+5)^2}{4}+(y-7)^2+4\]

OpenStudy (anonymous):

\[\frac{(x+5)^2}{16}+\frac{(y-7)^2}{4}=1\] is what i am trying to write, dang

OpenStudy (anonymous):

center at \[(-5,7)\] and major axis is horizontal. (16 is bigger than 4) so add and subtract 4 from the x - value to get the vertices. that gives the answer

OpenStudy (anonymous):

cooper, do you understand all those steps?

OpenStudy (anonymous):

Yes....after several days of working! Ha!